A body of mass 1kg and 2kg located (1,2) & (-1,3) respectively,then find out the centre of mass points.

Grade:12th pass

3 Answers

Arun
6 years ago
Dear Anil
 

body A  has mass of 1 kg  and location (1,2)

body B has mass of 2kg and location (-1,3)

mcxc = m1x1 + m2x2

(1+2) xc = (1*1) + (2*-1)

 xc  = -1/3

similarly 

mc yc = m1 y1 + m2 y2

(1=2) yc = (1*2)+(2*3)

 yc= 8/3

hence the co-ordinates of the center of mass are ( -1/3 , 8/3)

 

Regards

Arun (askIITians forum expert)

ANIL KUMAR PARIDA
6 years ago
Given that 2 bodies 1kg & 2kg located at (1,2) & (-1,3) respectively. Let M1=1kg , M2=2kg X1=1 , X2=-1 , Y1=2 , Y2= 3We know from cartesian coordinate system that centre of mass for X co-ordinate isXcm= M1*X1+M2*X2/M1+M2 =1*1+2(-1)/3 =1-2/3=-1/3Similarly for Y co-ordinate,Ycm=M1*X1+M2*X2/M1+M2 =1*2+2*3/3 =2+6/3 =8/3Hence the centre of mass points are (-1/3) & (8/3)
Chandan
6 years ago
take the second coordia